OpenStudy (anonymous):

in the interval (0,2pi) solve for x: cos^2x=cosx I NEED TO KNOW HOW TO DO THIS!

OpenStudy (misty1212):

\[z^2=z\\ z^2-z=0\\ z(x-1)=0\\ z=0,z=1\]

OpenStudy (danjs):

cos(x) = 1, When?

OpenStudy (danjs):

It is the x value on the unit circle,

OpenStudy (misty1212):

\[z=\cos(x)\] here and then \[\cos(x)=0\]\[\cos(x)=1\]

OpenStudy (jhannybean):

Then take the inverse to find x (technically the angle) \[x=\cos^{-1}(0)\]\[x=\cos^{-1}(1)\]

OpenStudy (danjs):



OpenStudy (danjs):

when is the x value either 0, or 1?

OpenStudy (danjs):

theta =

OpenStudy (jhannybean):